what is the momentum of a 40 000 ton (40 000 000 kg) oil tanker travelling at 5m/s?

1
Expert's answer
2021-11-03T06:22:41-0400
"p=mv=4\times10^7\ kg\times5\ \dfrac{m}{s}=2.0\times10^8\ \dfrac{kg\times m}{s}."